Consider an air-borne humanitarian mission, where food packages where dropped
with zero initial speed, from a helicopter hovering a height of h above the Earth
surface. Treating the package as particle of mass m, derive the horizontal deflection
from the particle caused by the Coriolis force acting on a particle falling freely in Earth’s
gravity field.
1
Expert's answer
2019-12-02T09:53:34-0500

The Coriolis force is much smaller in magnitude that the force of gravity.

In this case, Ώ is a small parameter.

We can writer

"z=h-\frac {g×t×t}{2} (1)"

The particle hits the ground when

"t=\sqrt {\frac {2h}{g}} (2)"

The net eastward deflection of the particle as strikes the ground is

"d=\frac {Ώ}{3} × \cosλ×\frac {8×h×h×h}{g}"